ca) Dre he Mohr cincle cleery correspond ns to drawns b) Determine panes por srons 2 11). Maximum sheer strin and correspondiCan you show me how to draw Mohr circle in detail scale and how to determine principal strains please? Also how to sketch the part (c) please? Please show me all part (a),(b) and (c) step by step. Thank you very much.
Answer #1

(c) The maximum shear strain is the radius of the mohr circle =3.687cm (y axis)

Multiply with the y axis scale you will get max. shear strain = 3.687*100*10-6 =368.7*10-6